# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v3

@@ Abstract
Deep supervised learning models require high volume of labeled data to attain sufficiently good results. Although, the practice of gathering and annotating such big data is costly and laborious. Recently, the application of self supervised learning (SSL) in vision tasks has gained significant attention. The intuition behind SSL is to exploit the synchronous relationships within the data as a form of self-supervision, which can be versatile. In the current big data era, most of the data is unlabeled, and the success of SSL thus relies in finding ways to utilize this vast amount of unlabeled data available. Thus it is better for deep learning algorithms to reduce reliance on human supervision and instead focus on self-supervision based on the inherent relationships within the data. With the advent of ViTs, which have achieved remarkable results in computer vision, it is crucial to explore and understand the various SSL mechanisms employed for training these models specifically in scenarios where there is limited labelled data available. In this survey, we develop a comprehensive taxonomy of systematically classifying the SSL techniques based upon their representations and pre-training tasks being applied. Additionally, we discuss the motivations behind SSL, review popular pre-training tasks, and highlight the challenges and advancements in this field. Furthermore, we present a comparative analysis of different SSL methods, evaluate their strengths and limitations, and identify potential avenues for future research.
@@
Self-supervised, Vision Transformer, Survey, Transformer
@@ 1 Introduction
Deep learning based algorithms have exhibited impressive results across various disciplines specially in computer vision (CV) [1] and natural language processing (NLP) [2]. Deep learning based models utilize a pre-training task on large datasets to enhance their performance. This primary step is often utilized as an initialization point and then the model is optimized for any specific use case. Self-supervised learning (SSL) is among the pre-training strategies for deep learning algorithms. Overall, SSL approach is driven by two major motivations. Firstly, a network that has been trained on an extensive data has already learned distinctive patterns that can be applied to subsequent tasks. This also helps to decrease the overfitting issue during training step. Secondly, parameters learned from extensive data provide effective parameter initialization, which enables faster convergence across different applications [3].
Although there is abundance of unlabeled web data in the era of big data, obtaining high-quality labeled data with human annotations can be costly. For instance, data labeling companies like Scale.com [4] charge \$6.4 per image for image segmentation labeling. Creating an image segmentation dataset with over 1 million high-quality samples (JFT-300M) could cost up to a million dollars. Usually, this process is time-intensive and inefficient [5]. Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations.
To tackle above-mentioned challenges in supervised learning, several learning strategies are presented, such as active learning [6], semi-supervised learning [7], and SSL [8]. Recently, Transformers [9] have emerged as an effective Neural Network (NN) architecture for leveraging deeper insights in the input data, specifically, the ViTs [10] are considered to be the alternative of CNNs [11]. ViTs have been successfully employed in tasks related to computer vision, including object detection [12], recognition [13] and semantic segmentation [14]. Notably, ViTs have demonstrated superior performance in image classification, when trained with large volume datasets such as JFT-300M [15].
In essence, by designing pretext tasks, a pre-training phase of network that encourage the model to learn significant patterns in the data itself, SSL methods enable ViTs to understand underlying structures of data without any supervision. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to employ the vast amount of unlabeled data that is readily available, allowing them to develop more robust and generalizable representations. SSL is beneficial in situations where labeled data is scarce or costly to acquire. Secondly, SSL provides a pathway for pre-training ViTs on larger volume of datasets, followed by of fine-tuning on downstream tasks, leading to improved performance and faster convergence [5].
@@ 1.1 Survey Scope and Contributions
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems [16], graphs [17], [18], sequential transfer learning [19], videos [20] and algorithms [5]. Vision Transformers (ViTs) are a prominent area of research in computer vision and have recently demonstrated several breakthroughs. However, the training of ViTs remains challenging. Therefore, this survey, contrary to existing surveys, explores the different SSL methods proposed for ViTs.
@@ 1.2 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. The reported SSL methods are segmented into 5 classes according to their unique features. The paper discuss the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them based on how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way of making use of input features. Frequently used abbreviations are listed in Table 1. The structure of paper is depicted in Fig 1. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceeding to section 2 which covers the advancements in SSL variants, while section 3 provide a taxonomy of the recent SSL architectures, respectively.
Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision. Section 5 presents various metrics and benchmarks for evaluating the effectiveness of SSL, while Section 6 provides a comparative analysis of state-of-the-art SSL mechanisms employing Vision Transformers. Section 7 highlights some open challenges in the domain of SSL, and Section 8 discusses different applications and future directions for SSL in Vision Transformers, including current challenges and potential developments. Finally, Section 9 concludes this survey.
Figure 1: Structure of various sections of the survey paper.
@@
GCMAE Global Contrast Masked Autoencode
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture and it is presented in seminal paper ”Attention Is All You Need” [9]. Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn representations from large amount of unlabeled data. ViTs need vast amounts of data to learn effective representations, and SSL provides a pathway for pre-training on large-scale datasets. By pre-training on large-scale datasets, ViTs can learn robust representations.
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learn the underlying relationship between its inputs. The framework of Energy-Based Models (EBMs) [32] can effectively describe this objective, where the aim of assigning high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pre-training [33], where layers of a deep neural networks are trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets [34]. Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
Contrastive learning became significant approach in self-supervised learning during the era of convolutional neural networks (CNNs) [21, 22]. The roots of contrastive learning can be found in the early work on metric learning and the development of Siamese networks [40] in the 1990s and early 2000s. Siamese Network architecture, introduced by Bromley et al. in 1993 [40], involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) [41] was a significant step forward.
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy (2016) [42] used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu (2018). [43] refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord (2018). [44] introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen (2020) [22] introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength, and in 2020, Caron (2020) [27] introduced SwAV, which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
The idea of self-distillation give a new dimension to SSL. Grill (2020) [28] boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict each other’s outputs. He et al. [21] introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron [45]) presented DINO, which uses self-distillation with no labels, leveraging VITs for SSL.
@@ 3 SSL based Techniques for ViTs
Figure 4: SSL Taxonomy developed using categorization of the approaches for self-supervised pre-training in ViTs.
3.0.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [21, 22]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks.
Literature based Review of Contrastive Learning
Wang and co-workers in 2022 [64] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@ 3.0.2 Knowledge Distillation Methods
Knowledge distillation transfers knowledge from a complex teacher model to a simpler student model without labeled data. This section explores literature on knowledge distillation in SSL applied to ViTs.
Literature based Review of Knowledge Distillation
Caron et al. [45] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
@@ 3.0.3 Clustering Methods
Su et al. (2023) [70] introduced a clustering-based SSL framework tailored for ViTs, aiming to enhance performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view which considers over an entire dataset. FLSL utilizes a bi-level feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally.
@@ 3.0.4 Generative Methods
Generative approaches, for example Generative Adversarial Networks (GANs) or Autoencoders, provide a promising avenue for pre-training ViTs. These models not only generate realistic images or reconstruct input images but also encode meaningful representations of visual content, facilitating enhanced learning of features in subsequent tasks.
@@
Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 [72] model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy on CIFAR-10, surpassing supervised Wide ResNet.
Bao et al. (2022) [73] introduced BEIT, that is Bidirectional Encoder representation from Image Transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual Transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
He et al., (2022) [53] proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks on the visible parts of image, and a small decoder rebuilds the original image usingthe learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pre-training was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.

### v4

@@ Abstract
Vision Transformers (ViTs) have recently demonstrated remarkable performance in computer vision tasks. However, their parameter-intensive nature and reliance on large amounts of data for effective performance have shifted the focus from traditional human-annotated labels to unsupervised learning and pretraining strategies that uncover hidden structures within the data. In response to this challenge, self-supervised learning (SSL) has emerged as a promising paradigm. SSL leverages inherent relationships within the data itself as a form of supervision, eliminating the need for manual labeling and offering a more scalable and resource-efficient alternative for model training. Given these advantages, it is imperative to explore the integration of SSL techniques with ViTs, particularly in scenarios with limited labeled data. Inspired by this evolving trend, this survey aims to systematically review SSL mechanisms tailored for ViTs. We propose a comprehensive taxonomy to classify SSL techniques based on their representations and pre-training tasks. Additionally, we discuss the motivations behind SSL, review prominent pre-training tasks, and highlight advancements and challenges in this field. Furthermore, we conduct a comparative analysis of various SSL methods designed for ViTs, evaluating their strengths, limitations, and applicability to different scenarios.
@@
Self-supervised Learning, Vision Transformer, Pre-training, Transfer Learning, Contrastive Learning, Generative Models
@@ 1 Introduction
Deep learning-based algorithms have achieved remarkable success across various disciplines, particularly in computer vision (CV) [1] and natural language processing (NLP) [2]. These models often utilize a pretraining task on large datasets to improve performance. This initial step typically serves as an initialization point, after which the model is fine-tuned for specific downstream tasks. Among pre-training strategies for deep learning, self-supervised learning (SSL) has emerged as a prominent approach. SSL approach is driven by two key motivations. Firstly, network trained on extensive datasets learn transferable patterns that reduce overfitting during training. Secondly, parameters derived from large-scale datasets provide effective initialization, enabling faster convergence between different applications [3].
Despite the abundance of unlabeled web data in the era of big data, developing high-quality labeled datasets through human annotation remains expensive. For instance, platforms such as Scale.com [4] charge approximately \$6.4 per image for the labeling task such as image segmentation. Building an image segmentation dataset containing over a million high-quality samples, like JFT-300M, can cost millions of dollars. Usually, this process is time-intensive and inefficient [5]. Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations.
To address these challenges, strategies like active learning [6], semi-supervised learning [7], and SSL [8] have gained traction. Recently, transformers [9] have emerged as an effective neural network (NN) architecture to leverage deeper insights into input data; specifically, ViTs [10] are considered the alternative to CNNs [11]. ViTs have been successfully employed in tasks related to computer vision, including object detection [12], recognition [13] and semantic segmentation [14]. In particular, ViTs have demonstrated superior performance in image classification, when trained with large-volume datasets such as JFT-300M [15].
In essence, SSL methods enable Vision Transformers (ViTs) to understand the underlying structures of data without supervision through pretext tasks, the self-designed objectives in a pre-training phase that encourage the model to learn significant patterns within the data itself. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to utilize large amounts of unlabeled data, fostering robust and generalizable feature representations. This approach is particularly advantageous when labeled data is limited or costly to produce. Secondly, SSL facilitates pre-training ViTs on extensive datasets, which can then be fine-tuned for specific downstream tasks. This process not only improves the performance of the model but also accelerates training convergence [5].
@@ 1.1 Survey Scope and Contributions
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems [16], graphs [17], [18], sequential transfer learning [19], videos [20] and algorithms [5]. Vision Transformers (ViTs) are a prominent area of research in computer vision and have recently demonstrated several breakthroughs. However, the training of ViTs remains a challenge. Therefore, in contrast to existing surveys, this paper explores the different SSL methods proposed for ViTs.
@@ 1.2 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. Then we discussed the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them according to how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way to make use of input features. Frequently used abbreviations are listed in Table. 1. The structure of paper is depicted in Fig. LABEL:fig:intro. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceed to section 2 which covers the advancements in SSL variants, while section 3 provides a taxonomy of the recent SSL architectures, respectively.
Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision. Section 5 presents various metrics and benchmarks to evaluate the effectiveness of SSL, while Section 6 provides a comparative analysis of state-of-the-art SSL mechanisms that employ Vision Transformers. Section 7 highlights some open challenges in the domain of SSL, and Section 8 discusses different applications and future directions for SSL in Vision Transformers, including current challenges and potential developments. Finally, Section 9 concludes this survey.
@@
1-2 GCMAE Global Contrast Masked Autoencode
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture which is presented in seminal paper ”Attention Is All You Need” [9]. Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn and understand feature representations from extensive unlabeled data. ViTs rely on extensive datasets to learn meaningful and effective representations while SSL offers a solution by enabling pretraining on large-scale datasets, allowing ViTs to acquire robust and transferable feature representations for downstream tasks.
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learn the underlying relationship between its inputs. The framework of Energy-Based Models (EBMs) [32] can effectively describe this objective, where the aim is to assign high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pretraining [33], where layers of a deep neural network is trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets [34]. Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
Contrastive learning became a significant approach in self-supervised learning during the era of CNNs [21, 22]. The roots of contrastive learning can be found in early work on metric learning and the development of Siamese networks [40] in the 1990s and early 2000s. Siamese Network architecture, introduced by Bromley et al. in 1993 [40], involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) [41] was a significant step forward.
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy (2016) [42] used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu (2018). [43] refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord (2018). [44] introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen (2020) [22] introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength and in 2020, Caron (2020) [27] introduced SwAV, which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
The idea of self-distillation gives a new dimension to SSL. Grill (2020) [28] boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict the outputs of each other. He et al. [21] introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron [45]) presented DINO, which uses self-distillation with no labels, leveraging VITs for SSL.
@@ 3 SSL based Techniques for ViTs
3.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [64, 21, 22]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks.
Literature based Review of Contrastive Learning Li et al. [65] suggested a methodology to enhance domain generalization in mammographic image analysis using contrastive learning. Their approach consists of two stages: first, a multi-style and multi-view contrastive learning pretrainig phase, where CycleGAN and image blending techniques generate diverse mammogram styles to enhance feature robustness; second, a fine-tuning stage for downstream tasks such as mass detection, mass matching, BI-RADS classification, and breast density classification. By learning style-variant representations, their method demonstrated improved performance across multiple mammographic datasets, outperforming existing domain generalization techniques. Wang and co-workers in 2022 [66] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@ 3.1.1 Knowledge Distillation Methods
Knowledge distillation transfers knowledge from a complex teacher model to a simpler student model without labeled data. This section explores literature on knowledge distillation in SSL applied to ViTs. Literature based Review of Knowledge Distillation Caron et al. [45] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
@@ 3.1.2 Clustering Methods
Hemalatha et al. (2025) [72] introduced a novel approach to enhance cervical cancer detection by using self-supervised learning techniques on diverse cell images.
Su et al. (2023) [73] introduced a clustering-based SSL framework tailored for ViTs, aiming to improve the performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view, which considers over an entire dataset. FLSL utilizes a bilevel feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally.
@@ 3.1.3 Generative Methods
Generative approaches, for example Generative Adversarial Networks (GANs) or Autoencoders, provide a promising avenue for pre-training ViTs. These models not only generate realistic images or reconstruct input images, but also encode meaningful representations of visual content, facilitating enhanced learning of features in subsequent tasks.
@@
Heba et al. (2025) [64] suggest that both contrastive learning and in-painting are effective auxiliary tasks for SSL in the medical domain, leading to a 5.26% increase in accuracy for polyp classification compared to other weight initialization methods.
Jie et al. (2025) [75] proposed a self-supervised BM-GAN to learn morphological changes in MRI over the course of a disease. BM-GAN, learns both forward and backward temporal dependencies to interpolate missing time-series data at arbitrary time points. By integrating self-supervised learning and a bidirectional GAN framework, the model ensures realistic reconstruction without labeled data, offering a robust solution for handling irregular and incomplete time-series datasets.
Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 [76] model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy in CIFAR-10, surpassing supervised wide ResNet.
Bao et al. (2022) [77] introduced BEIT, that is, bidirectional encoder representation from image transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
He et al. (2022) [53] proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks at the visible parts of the image, and a small decoder rebuilds the original image using the learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pre-training was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.